Accounting - Chapter 10

A. Legacy issues $630,000 of 9.0%, four-year bonds dated January 1, 2019, that pay interest semiannually on June 30 and December 31. They are issued at $571,310 when the market rate is 12%.

a)Prepare an effective interest amortization table for the bonds' first two years

b) Prepare the journal entries to record the first interest payments
Answer #1

a) 

Semiannual period-end

Unamortized   Discount

Carrying   Value

01/01/2021

58690

571310

06/30/2021

51354

578646

12/31/2021

44018

585982

06/30/2022

36681

593319

12/31/2022

29345

600655

Total discount= 630000-571310= 58690

ð  Per semiannual period = 58690/8= 7336.25 ( 7336)

Interest expense = (630000*9%*1/2)= 28350

Total Bond Interest expense = (630000*9%*1/2) *8 + 630000 – 571310= 285490

Bond Interest expense per semiannual period end = 285490/8= 35686.25

b)

no

Date

General journal

debit

credit

1

June 30

Bond interest expense

35686




Discount on bonds payable


7336



Cash


28350


answered by: Lovestudying